News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 7 July 2000 12.30pm EDT

During the last week, Chandra completed the observing schedule as planned.

ACIS focal plane temperature was successfully set to -110 deg C for multiple calibration observations of two supernova remnants on July 3, 5, and 6. This completed the set of calibration data needed for analysis of observations prior to 31 Jan 2000. The focal plane was cooled back down to the normal -120 deg C setting for science targets.

A routine real-time procedure was carried out to dump the CPE warm start counter, which showed the correct value of (1), as expected.

Regression testing of OFLS 9.0 is proceeding. Current work is on the Command Management and the Mission Planning and Scheduling systems.

The planned observations for the coming week are as follows.

molecular cloud MBM 12 ACIS-S          Jul  9
Wolf-Rayet Star WR 147 HRC-I           Jul 10
Low Mass Star VB 8     ACIS-S
1WGA J1340.1+2743      ACIS-S
SuperNova 1006-NEast   ACIS-S
Radiation Belt                         Jul 11
M33                    ACIS-I          Jul 12
NGC 1333               ACIS-I
M33                    ACIS-S          Jul 13
Comet c/LINEAR 1999    ACIS-S          Jul 14 (8 pointings)
Radiation Belt
Comet c/LINEAR 1999    ACIS-S          Jul 15 (8 pointings)
molecular cloud MBM 12 ACIS-S
MITG J0414+0534        ACIS-S
Supernova Rem. 3C 58   ACIS-S


All spacecraft subsystems continued to operate nominally.

Return to Status Reports